# Sourced Carrollian Fluids Dual to Black Hole Horizons

### v1

@@ I Introduction
The study of the Carrollian limit began with the works of Lévy-Leblond [23] and Gupta [24] as a mathematical exploration of the contraction [25] of the Poincaré group, in which the speed of light vanishes. These studies evolved into an active field in physics in the last decade [26]. Some of the works on the field focused on the relations between the asymptotic symmetries of asymptotically flat spacetimes with Carroll groups and their use in flat holography [27, 28, 29, 30]; Carrollian fluids as the limit of relativistic hydrodynamics [31, 32, 33, 34]; the interacting particle motion [35] and dynamics of extended objects [36]; Carrollian limits of relativistic field theories [37, 38, 39, 40, 41, 42, 43, 44, 45] [^1] Note also the early work (1998) by Dautcourt [46] on the “ultrarelativistic” limit of general relativity.; applications in cosmology with emphasis on dark energy [47, 48]. Here, we focus on the special case of Carrollian fluids dual to black hole horizons and utilize our observations on the gravitational side of the duality to inform the Carrollian side.
@@ II.2 The Metric & Gauge Choices: Near Horizon
$n^{a}\partial_{a}=\partial_{v}+{2\rho}\partial_{r}+V^{A}\partial_{A},$ $\left(e_{A}\right)_{a}\mathrm{d}x^{a}$ $=V_{A}\mathrm{d}v+q_{AB}\mathrm{d}x^{B}.$
@@ V Perturbative Treatment
$g_{ab}^{(0)}=-\frac{\tilde{r}}{2M}\mathrm{d}v^{2}+2\mathrm{d}v\,\mathrm{d}\tilde{r}+4M\left(M+\tilde{r}\right)\left(\mathrm{d}\theta^{2}+\sin{\theta}^{2}\mathrm{d}\phi^{2}\right).$ (41)
@@
    $\partial_{v}\Omega^{(1)}_{AB}=2M\partial_{v}\lambda_{AB}^{(1)}+\frac{1}{2}\lambda_{AB}^{(1)}-\frac{1}{2M}\Omega_{AB}^{(1)}-2M{\nabla}_{\langle A}\mathcal{H}_{B\rangle}^{(1)}$ (54)
@@
    $\qquad+4M\left({\nabla}_{\langle A}\mathcal{H}_{B\rangle}^{(2)}+\mathcal{H}^{(1)}_{\langle A}\mathcal{H}^{(1)}_{B\rangle}\right)+2M{\partial}_{A}\Phi^{(1)}{\partial}_{B}\Phi^{(1)}-M\Omega^{\text{(Sph)}}_{AB}\left(\Omega^{CD}_{\text{(Sph)}}{\partial}_{C}\Phi^{(1)}{\partial}_{D}\Phi^{(1)}\right)$
@@
    $\qquad-M\left({\nabla}_{A}\mathcal{H}^{A(2)}+\mathcal{H}^{(1)}_{A}\mathcal{H}^{A(1)}\right)+\frac{M}{2}\hat{R}^{(2)}-M\Omega^{AB}_{\text{(Sph)}}\partial_{A}\Phi^{(1)}\partial_{B}\Phi^{(1)}$
@@ V.1 Spin-weighted Functions
    $\partial_{v}C$ $=2M\partial_{v}S+\frac{1}{2}S-\frac{1}{2M}C+2k\left[Ms-c\right]$
    $\quad+4M\left(\eth H+h^{2}\right)+2M\left(\eth\varphi\right)^{2}$ (73)
@@ V.2.3 Horizon angular momentum
Contrarily, the angular momentum computed from the Hájíček field has a completely different behavior. The solid orange curve rises exponentially before $v=0$. This is due to the secular terms in Eq.s 73 & 74, which permeate to the angular momentum computed from Eq. 91. Therefore, at some $v$ around $v=0$, the perturbative approximation fails to capture the physical behavior.
@@ Acknowledgements.
We would like to thank Luca Ciambelli, Puttarak Jai-akson, Jaime Redondo-Yuste and Sizheng Ma for useful discussions. LL acknowledges support from the Natural Sciences and Engineering Research Council of Canada through a Discovery Grant. LL also thanks financial support via the Carlo Fidani Rainer Weiss Chair at Perimeter Institute and CIFAR. This research was supported in part by Perimeter Institute for Theoretical Physics. Research at Perimeter Institute is supported in part by the Government of Canada through the Department of Innovation, Science and Economic Development and by the Province of Ontario through the Ministry of Colleges and Universities.

### v2

@@ I Introduction
The study of the Carrollian limit began with the works of Lévy-Leblond [23] and Gupta [24] as a mathematical exploration of the contraction [25] of the Poincaré group, in which the speed of light vanishes. These studies evolved into an active field in physics in the last decade [26]. Some of the works on the field focused on the relations between the asymptotic symmetries of asymptotically flat spacetimes with Carroll groups and their use in flat holography [27, 28, 29, 30, 31]; Carrollian fluids as the limit of relativistic hydrodynamics [32, 33, 34, 35]; the interacting particle motion [36] and dynamics of extended objects [37]; Carrollian limits of relativistic field theories [38, 39, 40, 41, 42, 43, 44, 45, 46, 47]  [^1] Note also the early work (1998) by Dautcourt [48] on the “ultrarelativistic” limit of general relativity.; applications in cosmology with emphasis on dark energy [49, 50]. Thermodynamic properties of Carrollian fluids have also been studied in recent works: for example, in [51] from the point of view of symmetries (as was developed in a general framework in [52]), and in [53] within a holographic setup by relating the gravitational radiation with the dissipation in the dual Carrollian theory on the boundary. Here, we focus on the special case of Carrollian fluids dual to black hole horizons and utilize our observations on the gravitational side of the duality to inform the Carrollian side.
@@ II.2 The Metric & Gauge Choices: Near Horizon
$n^{a}\partial_{a}=\partial_{v}+{2\rho}\partial_{r}+V^{A}\partial_{A},$ $\hskip-5.0pt\left(e_{A}\right)_{a}\mathrm{d}x^{a}$ $=V_{A}\mathrm{d}v+q_{AB}\mathrm{d}x^{B}.$
@@ V Perturbative Treatment
$g_{ab}^{(0)}=-\frac{\tilde{r}}{2M}\mathrm{d}v^{2}$ $+2\mathrm{d}v\,\mathrm{d}\tilde{r}$
$+4M\left(M+\tilde{r}\right)\left(\mathrm{d}\theta^{2}+\sin{\theta}^{2}\mathrm{d}\phi^{2}\right).$ (41)
@@
    $\partial_{v}\Omega^{(1)}_{AB}$ $=2M\partial_{v}\lambda_{AB}^{(1)}+\frac{1}{2}\lambda_{AB}^{(1)}$
    $\quad-\frac{1}{2M}\Omega_{AB}^{(1)}-2M{\nabla}_{\langle A}\mathcal{H}_{B\rangle}^{(1)}$ (54)
@@
    $\qquad+4M\left({\nabla}_{\langle A}\mathcal{H}_{B\rangle}^{(2)}+\mathcal{H}^{(1)}_{\langle A}\mathcal{H}^{(1)}_{B\rangle}\right)+2M{\partial}_{A}\Phi^{(1)}{\partial}_{B}\Phi^{(1)}-M\Omega^{\text{(Sph)}}_{AB}\left(\Omega^{CD}_{\text{(Sph)}}{\partial}_{C}\Phi^{(1)}{\partial}_{D}\Phi^{(1)}\right)$ (58)
@@
    $\qquad-M\left({\nabla}_{A}\mathcal{H}^{A(2)}+\mathcal{H}^{(1)}_{A}\mathcal{H}^{A(1)}\right)+\frac{M}{2}\hat{R}^{(2)}-M\Omega^{AB}_{\text{(Sph)}}\partial_{A}\Phi^{(1)}\partial_{B}\Phi^{(1)}$ (60)
@@ V.1 Spin-weighted Functions
    $\partial_{v}C$ $=2M\partial_{v}S+\frac{1}{2}S-\frac{1}{2M}C+2k\left[Ms-c\right]+4M\left(\eth H+h^{2}\right)+2M\left(\eth\varphi\right)^{2}$ (75)
@@ V.2.3 Horizon angular momentum
Contrarily, the angular momentum computed from the Hájíček field has a completely different behavior. The solid orange curve rises exponentially before $v=0$. This is due to the secular terms in Eq.s 75 & 76, which permeate to the angular momentum computed from Eq. 93. Therefore, at some $v$ around $v=0$, the perturbative approximation fails to capture the physical behavior.
@@ Acknowledgements.
We would like to thank Luca Ciambelli, Puttarak Jai-akson, Jaime Redondo-Yuste, and Sizheng Ma for useful discussions. LL acknowledges support from the Natural Sciences and Engineering Research Council of Canada through a Discovery Grant. LL also thanks financial support via the Carlo Fidani Rainer Weiss Chair at Perimeter Institute and CIFAR. This research was supported in part by Perimeter Institute for Theoretical Physics. Research at Perimeter Institute is supported in part by the Government of Canada through the Department of Innovation, Science and Economic Development and by the Province of Ontario through the Ministry of Colleges and Universities.